# SiM3D: Single-instance Multiview Multimodal and Multisetup 3D Anomaly Detection Benchmark

### v1

@@ 6 Final Discussion
We introduced SiM3D, the first benchmark and dataset focused on integrating multiview and multimodal information for comprehensive 3D ADS. Our experiments highlight several open challenges. First, a need for methods that can process multiview inputs more effectively, e.g., by taking all views as input rather than integrating singleview outputs, and that can be trained with a very limited number of nominal samples, e.g., a single object instance. Second, training with synthetic data requires more robust techniques to address domain-shift challenges. Finally, the current literature lacks 3D backbones capable of processing high-resolution point clouds. Employing depth maps seems a promising strategy to facilitate this processing, yet a foundational model for depth maps is still missing. We believe that SiM3D may serve as a strong foundation for tackling these challenges and fostering future research toward 3D ADS.
@@
Figure D: Qualitatives. Selected views with defects (left) and Anomaly Volumes for several methods (right) downsampled for visualization.
@@
For the following baselines, we process greyscale images $\{I_{i}\}_{i=1}^{n}$, depth maps $\{D_{i}\}_{i=1}^{n}$ or organized point clouds $\{P_{i}\}_{i=1}^{n}$ downsampled to the common size of $1540\times 1540\times 3$, the highest achievable with the available hardware.
Input data processed with WideResNet101, either greyscale images or depth maps, returns features from the second and third layers, yielding feature maps with dimensions equal to $28\times 28\times 1536$. Input data processed with DINO-v2, either greyscale images or depth maps, return features from the last layer, yielding feature maps with dimensions equal to $110\times 110\times 768$. Input data processed with FPFH features, namely point clouds, are processed considering a voxel size equal to 2 mm in order to match the resolution attainable from the 3D voxel ground-truths and yield feature maps with dimensions equal to $1540\times 1540\times 33$. The spatial resolution of these maps is then downsampled to match either the spatial resolution of the WideResNet101 or DINO-v2 features, hence $28\times 28$ or $110\times 110$.
@@
We implement EfficientAD by disabling the Teacher normalisation since there is no validation set available, and we upsample the intermediate encoder outputs to match the size of the Teacher and the Students fed with $1540\times 1540\times 3$ images. Moreover, we train the Students for 1000 epochs, unlike the 70000 expected from the adopted implementation, since the loss tends to stall earlier due to the limited number of training images that characterise our single-instance setup.
@@
S.7 Attribution of existing assets
@@
Figure E: SiM3D dataset overview. From top to bottom: the single-instance real and synthetic training samples for objects Rubbish Bin and Wicker Vase, one of the anomalous samples from the test set.
Figure F: SiM3D dataset overview. From top to bottom: the single-instance real and synthetic training samples for objects Bathroom Furniture and Container, one of the anomalous samples from the test set.
Figure G: SiM3D dataset overview. From top to bottom: the single-instance real and synthetic training samples for objects Plastic Vase and Wooden Stool, one of the anomalous samples from the test set.
S.8 Ethical statement
@@
S.9 Additional qualitative results
We report in Fig. D additional qualitative results concerning the classes of SiM3D which have not been displayed in in Fig. 5.
S.10 Extended dataset visualizations
Akin to Fig. 1 of the main paper, in Fig. E, Fig. F and Fig. G, we show training samples, both real and synthetic, as well as defective test samples for other object types present in the SiM3D dataset.

### v2

@@ 6 Final Discussion
We introduced SiM3D, the first benchmark and dataset focused on integrating multiview and multimodal information for comprehensive 3D ADS. Our experiments highlight several open challenges. First, a need for methods that can process multiview inputs more effectively, e.g., by taking all views as input rather than integrating singleview outputs, and that can be trained with a very limited number of nominal samples, e.g., a single instance. Second, training with synthetic data requires more robust techniques to address domain-shift challenges. Finally, the current literature lacks 3D backbones capable of processing high-resolution point clouds. Employing depth maps seems a promising strategy to facilitate this processing, yet a foundational model for depth maps is still missing. We believe that SiM3D may serve as a strong foundation for tackling these challenges and fostering future research toward 3D ADS.
Acknowledgment
We acknowledge financial support from the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call No. 1409 published on 14 September 2022 by the Italian Ministry of University and Research (MUR), funded by the European Union.
This work was also partially supported by the FSE+ 2021–2027 program under Article 24, paragraph 3, letter a) of Law 240/2010 and its subsequent amendments, as well as by Regional Resolution No. 693/2023 (Project Code: 2023-20090/RER), Grant CUP: J19J23000730002.
We sincerely thank Lucia Gasperini for her contribution to dataset acquisition and labelling. We also extend our gratitude to Manuel Turrini for his support in setting up the acquisition process.
@@ Acknowledgment
Figure D: Qualitatives. Views with defects (left) and Anomaly Volumes for several methods (right) downsampled for visualisation.
Figure E: Multimodal methods qualitatives. Views with defects (left) and Anomaly Volumes for all multimodal methods (right) downsampled for visualisation.
@@
For the following baselines, we process greyscale images $\{I_{i}\}_{i=1}^{n}$, depth maps $\{D_{i}\}_{i=1}^{n}$ or organized point clouds $\{P_{i}\}_{i=1}^{n}$ downsampled to the common size of 1540×1540×31540$\times$ 1540$\times$ 31540 × 1540 × 3, the highest achievable with the available hardware.
Input data processed with WideResNet101, either greyscale images or depth maps, returns features from the second and third layers, yielding feature maps with dimensions equal to 28×28×153628$\times$ 28$\times$ 153628 × 28 × 1536. Input data processed with DINO-v2, either greyscale images or depth maps, return features from the last layer, yielding feature maps with dimensions equal to 110×110×768110$\times$ 110$\times$ 768110 × 110 × 768. Input data processed with FPFH features, namely point clouds, are processed considering a voxel size equal to 2 mm in order to match the resolution attainable from the 3D voxel ground-truths and yield feature maps with dimensions equal to 1540×1540×331540$\times$ 1540$\times$ 331540 × 1540 × 33. The spatial resolution of these maps is then downsampled to match either the spatial resolution of the WideResNet101 or DINO-v2 features, hence $\times 28$ 28 or $\times 110$ 110.
@@
We implement EfficientAD by disabling the Teacher normalisation since there is no validation set available, and we upsample the intermediate encoder outputs to match the size of the Teacher and the Students fed with 1540×1540×31540$\times$ 1540$\times$ 31540 × 1540 × 3 images. Moreover, we train the Students for 1000 epochs, unlike the 70000 expected from the adopted implementation, since the loss tends to stall earlier due to the limited number of training images that characterise our single-instance setup.
@@
Asymmetric Student-Teacher. AST [32] is a singleview, multimodal ADS method that employs EfficientNet-B5 to extract features from RGB images and depth maps. Such features are subsequently employed to optimise a Normalizing Flow as Teacher network and a feed-forward network as a Student network. The idea is that, after optimisation, both networks are able to reconstruct nominal samples, begetting low discrepancies, while failing to reconstruct anomalous samples. Since these two networks present different architectures, the way in which they fail to reconstruct anomalous samples is different, hence discrepancies will be exacerbated, highlighting anomalies. Due to the fact that this method work on the features, it is triviality extendable to multiple modalities.
Given that SiM3D contains high-resolution images and point clouds, we deployed AST by passing to the framework 1540×1540×31540$\times$ 1540$\times$ 31540 × 1540 × 3 images and depth maps.
Pl. Stool Rub. Bin W. Vase B. Furn. Cont. Pl. Vase W. Stool Sink Cab. Mean
FID real2test 10.088 8.844 7.345 15.784 12.590 9.560 8.910 23.750 12.108
FID synth2test 48.744 44.565 57.302 40.900 51.900 55.194 24.040 34.060 44.588
$\Delta$ FID 38.656 35.721 49.957 25.116 39.310 45.634 15.130 10.310 32.480
$\Delta$ I-AUROC (PatchCore) 0.458 0.061 0.223 0.158 0.371 0.121 0.393 0.476 0.282
$\Delta$ V-AUPRO (PatchCore) 0.044 0.012 0.003 0.229 0.061 0.019 0.114 0.115 0.075
Table C: Data quality assessment and impact on performance. The average Fréchet Inception Distance between the train samples and the test samples for both real2real and synth2real setups is reported.
Figure F: Outliers. The same test views inferred with the same model, trained either with real2real or synth2real train set, highlights the presence of outliers in the synth2real scenario.
S.7 Assessment of data generation quality and insight of failure cases of the synth2real setup
We report in Tab. C the Fréchet Inception Distance (FID) between the train set, either from the real2real or synth2real setup, and the test set, along with the performance gaps ($\Delta$ I-AUROC, $\Delta$ V-AUPRO) for PatchCore. We do not observe any correlation between $\Delta$ FID and $\Delta$ I-AUROC or $\Delta$ V-AUPRO. Instead, in the synth2real scenario, we noticed strong outliers in anomaly maps of nominal samples (see Fig. F), for some objects. These outliers have a strong impact on detection since they increase the false detection rate, while affecting the segmentation performance much less.
S.8 Ancillary experiments to assess the SiM3D challenges
As suggested during the peer-reviewing process, we ran the following experiments to gain more insight into the unique challenges set forth by SiM3D:
  • We treated the task as single-view ADS by deploying PatchCore, obtaining a mean I-AUROC of 0.488 vs. 0.754 of our multiview approach, highlighting the necessity of addressing the task in a multiview fashion. Notice that we cannot compare the segmentation performance since the ground-truths are voxel grids;
  • We ran the official implementation of PatchCore on grayscale MVTec AD, obtaining a mean I-AUROC of 0.988 vs. 0.991, and a mean AUPRO@30% of 0.929 vs. 0.935, both official results from [30], highlighting that benchmark complexity does not stem from grayscale images.
As for metrics, I-AUROC is directly comparable to other benchmarks; hence, the general lower performance confirms that SiM3D is more challenging. Segmentation performance cannot be compared since the ground-truths of SiM3D are voxel grids.
S.9 Assessment of baselines on 2D and multimodal anomalies
As suggested during the peer-reviewing process, we performed a comparative analysis of some baselines while disentangling the different kinds of anomalies. To this end, starting from Tab. B, we split the anomalies into 2D and 3D+multimodal anomalies, hence, unifying 3D defects with the multimodal ones. Indeed, we argue that, while it is possible to unambiguously consider some anomalies, such as marker strokes and white-outs, as 2D-only, it is more difficult to pinpoint 3D-only anomalies, as structural deviations in the geometry, more often than not, tend to manifest themselves also in the image space.
After that, we selected the best-performing algorithm working on images (PatchCore w/ DINO-v2) and multimodal data (AST), The experiments, reported in Tab. D, show that PatchCore, the image-based method, tend to perform best on 2D anomalies and worse on multimodal anomalies, obtaining an overall weaker performance when all the anomalies are considered. On the other hand, AST, the multimodal method, tend to perform well on both 2D and multimodal anomalies, obtaining an overall stronger performance when all the anomalies are considered. These findings further confirm the necessity of a multimodal analysis when working in the SiM3D setup.
Method Modality Anomalies Detection Segmentation
PatchCore w/ DINO-v2 RGB All 0.500 0.958 0.636 0.622 0.578 0.563 1.000 0.563 0.678 0.745 0.469 0.775 0.792 0.709 0.753 0.435 0.690 0.671
PatchCore w/ DINO-v2 RGB 2D 0.646 0.902 0.602 0.723 0.678 0.498 0.980 0.612 0.705 0.832 0.498 0.674 0.801 0.718 0.687 0.455 0.678 0.667
PatchCore w/ DINO-v2 RGB Multimodal 0.425 0.573 0.534 0.439 0.592 0.632 0.754 0.587 0.567 0.276 0.354 0.456 0.548 0.698 0.603 0.404 0.596 0.491
AST w/ EffNet-B5 RGB + Depth All 0.950 0.927 0.785 0.474 0.542 0.470 0.428 0.925 0.687 0.750 0.503 0.792 0.807 0.716 0.764 0.467 0.798 0.699
AST w/ EffNet-B5 RGB + Depth 2D 0.910 0.905 0.773 0.427 0.539 0.320 0.423 0.892 0.648 0.654 0.443 0.730 0.687 0.708 0.797 0.303 0.679 0.625
AST w/ EffNet-B5 RGB + Depth Multimodal 0.945 0.895 0.698 0.410 0.498 0.543 0.413 0.904 0.663 0.738 0.475 0.765 0.789 0.893 0.683 0.564 0.723 0.703
Table D: Anomaly detection and segmentation results considering different kinds of anomalies.
S.10 Attribution of existing assets
@@
Figure G: SiM3D dataset overview. From top to bottom: the single-instance real and synthetic training samples for objects Rubbish Bin and Wicker Vase, one of the anomalous samples from the test set.
Figure H: SiM3D dataset overview. From top to bottom: the single-instance real and synthetic training samples for objects Bathroom Furniture and Container, one of the anomalous samples from the test set.
Figure I: SiM3D dataset overview. From top to bottom: the single-instance real and synthetic training samples for objects Plastic Vase and Wooden Stool, one of the anomalous samples from the test set.
S.11 Ethical statement
@@
S.12 Additional qualitative results
We report in Fig. D additional qualitative results concerning the classes of SiM3D which have not been displayed in in Fig. 5. We additionally report in Fig. E qualitative results for top-performer multimodal methods reported in Tab. 5.
S.13 Extended dataset visualizations
Akin to Fig. 1 of the main paper, in Fig. G, Fig. H, and Fig. I, we show training samples, both real and synthetic, as well as defective test samples for other object types present in the SiM3D dataset.